OpenStudy (rippedblue42):

If p(x)/q(x)= x+4 with a remanded of -11 and Q(x)=x+1, find p(x)

OpenStudy (mathmale):

\[\frac{ p(x) }{ q(x) }=x+4\] If you want to find p(x), multiply both sides of this equation by q(x)=x+1. Then add 11 to your result.

OpenStudy (sshayer):

p(x)=(x+4)q(x)-11 p(x)=(x+4)(x+1)-11

OpenStudy (mathmale):

Note that I suggested ADDING 11 to y our result, whereas sshayer suggested SUBTRACTING 11. You can determine which is correct by dividing p(x) by q(x)=x+1.